OpenStudy (rock_mit182):

The sum of two numbers is 40. the smaller number is 20 less than the larger number. what are the numbers ?

OpenStudy (anonymous):

you have to make an equation for this

OpenStudy (anonymous):

Let small number = x large number = y

OpenStudy (anonymous):

name the larger number x

OpenStudy (anonymous):

and try to express the smaller number using the larger number (x)

OpenStudy (rock_mit182):

x = smaller number y = larger number 1) x+y = 40 2) x = y - 20 or x-y = -20 system of equations: x + y = 40 x - y =-20 solving: 2x = 20 --> x = 10.

OpenStudy (anonymous):

no you have to use one variable

OpenStudy (anonymous):

So from first statement sum of them = x + y = 40

OpenStudy (rock_mit182):

replacing in 1) 10 + y = 40 --> y = 30

OpenStudy (anonymous):

x:large number

OpenStudy (rock_mit182):

\[x=10,y =30\] am i right gusy, I'm just checking

OpenStudy (rock_mit182):

hey please someone. is my work right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Absolutely right x =10 y = 30

OpenStudy (anonymous):

but you could solve it with the following way to: let x:large number and x-20:small number so: x-20+x=40 solve it for x ....

OpenStudy (rock_mit182):

Ok thanks to both of you !

OpenStudy (rock_mit182):

i know both diserve the medal

OpenStudy (rock_mit182):

so the fair thing to do would be medals for everyone